# Enhancing the accuracy of observable distributions for galaxies classified in the Projected Phase Space Diagram

### v1

@@ 3.2 Applying the method
    Interlopers: As said before, for this class, our method provides no improvement at all, but this should not be considered as a problem since there are virtually no differences between the intrinsic, predicted and recovered colour distributions, as in the previous case. In all five mass bins we obtain $S_{\rm ITL}^{\rm rec}(M)>S_{\rm ITL}^{\rm pred}(M)$, but note that all numerical values ($\sim 10^{-3}$) are an order of magnitude smaller than for IN.
3.3 The effects of a mass dependence of the matrix $\mathbf{C}$
We have explored the effect on our method of a possible dependence of the confusion matrix on the mass of galaxies. For each mass bin considered in Fig. 2, we compute a confusion matrix, $\mathbf{C}(M)$, using only those galaxies in the test sample that have stellar mass in the bin. The matrices so obtained (not shown here) do have a mild dependence on mass. However, this dependence do not change the results discussed in the previous subsection in a significant way. In Fig. 3, we present the resulting values of $S_{i}^{\rm rec}(M)$ as orange lines, where it is evident that the differences compared to the case of a unique $\mathbf{C}$ are really small. We can safely affirm there is no need to use a mass (or absolute magnitude) dependent confusion matrix. Furthermore, a mass dependent confusion matrix implies a greater dependence on the specifics of how the galaxy evolution model works. It is desirable to minimise reliance on the model as much as possible.
@@ 4 Application to an observed sample of X-ray clusters of galaxies
As an example of the potential of the proposed statistical technique to correct the effects of misclassification, we apply our method to the sample of X-ray clusters and SDSS galaxies used in Martínez et al. (2023). The sample comprises 104 X-ray clusters drawn from the works by Coenda & Muriel (2009) and Muriel & Coenda (2014), where the authors compute a number of properties of clusters originally catalogued by Popesso et al. (2004) and Bohringer et al. (2000), respectively. Martínez et al. (2023) searched for all spectroscopic SDSS-DR7 (Abazajian et al. 2009) galaxies within cylindrical volumes in redshift space centred in the clusters, and with radius $3\times R_{200}$ in projection and a longitude of ±$\pm 3\times\sigma$ alongside the line-of-sight of the cluster. Galaxies were classified in the five classes under study using their positions in the PPSD, the roger computed probabilities, and the thresholds of Coenda et al. (2022).
Using the confusion matrix computed in Sect. 3.2, we estimate the predicted and recovered distributions of the (${}^{0.1}(u-r)$ colour of the five classes in four bins in stellar mass. The resulting distributions are shown in Fig. 4. The outcome of applying our method to this sample of observed galaxies and clusters qualitatively agrees with our findings in Sect. 3.2, which were derived from applying the method to a simulated galaxy catalogue. We can observe basically the same trends of Fig. 3. For cluster members, the distributions recovered by our method eliminate most of the contribution from blue galaxies, this is more notorious at the lower mass bins. According to the insights of Sect. 3.2, these blue galaxies are mostly actual recent infallers. In the case of backsplash galaxies, the main feature we obtain is the removal of the contamination from blue galaxies, particularly at the lowest mass bin. According to the confusion matrix we are using (see Sect. 3.2), some $70\%$ of the contamination should be due to erroneously classified infalling galaxies. Regarding recent infallers, the result of our method is most notably seen in the removal of a population of red, almost exclusively cluster members, most clearly at the lower mass bins. As in Sect. 3.2, for infalling galaxies, the method gives results very similar to the raw predictions. Finally, for interlopers, once again, there are virtually no differences between the predicted and recovered colour distributions.
@@ 5 Conclusions
We propose a method to recover the distributions of observed properties of galaxies classified according to their position in the projected phase space diagram. The method involves the knowledge of the confusion matrix, or at least an estimation of it, which is inverted and applied to observed distributions. We have tested this method on simulated clusters and galaxies and applied it to a sample of galaxies and X-ray clusters from the Sloan Digital Sky Survey. For both, simulated and observed galaxies, we classify galaxies in the PPSD using the code roger of de los Rios et al. (2021), plus the criteria by Coenda et al. (2022). In particular, we have utilised colour as the parameter for which we analyse the distributions. These two choices, namely, the classification using roger plus thresholds, and colour, are not essential to our objectives. Any classification scheme over the PPSD, such as those by Rhee et al. (2017) and Pasquali et al. (2019), and any galaxy property, would be suitable provided one has an estimation of the confusion matrix.
Using the inversion of the confusion matrix along with our roger-based classification provides better estimation of the distribution of colours of galaxies in and around clusters. The method is successful at decontaminating different classes of galaxies from the contamination of other classes, and it is particularly useful over those regions of the PPSD where the overlapping between classes complicate the classification most. The method proposed allows for further analyses on galaxies when the classification is an issue. It is straightforward to reconstruct some statistics of galaxy populations as a function of stellar mass (or absolute magnitude). For instance, from Fig. 4, we could compute the median of the colour distribution as a function of mass for the blue and red galaxy populations of each galaxy class.
Regarding the results we obtain for the five classes, there are some main conclusions we can draw. The bulk of galaxies in clusters that are blue and less massive than $M\simeq 10^{10.5}h^{-1}M_{\odot}$ are mainly recent infallers. In terms of colour, backsplash galaxies are on average intermediate between cluster members and recent infallers, while being distinctly redder than infalling galaxies. Thus, a single excursion through the inner regions of a cluster produces significant changes in galaxy colour, due to a quench in star formation (e.g. Hough et al. 2023; Ruiz et al. 2023).
@@ Appendix A Additional figure
Figure 4: Application of our method to the sample of galaxies in and around X-ray clusters used by Martínez et al. (2023). As in Fig.2, rows consider different classes and columns different stellar mass ranges. Green shaded histograms: predicted colour distributions. Violet shaded histograms: recovered colour distributions.

### v2

@@ 3.2 Applying the method
    Interlopers: As said before, for this class, our method provides no improvement at all, but this should not be considered as a problem since there are virtually no differences between the intrinsic, predicted and recovered colour distributions, as in the previous case. In all five mass bins we obtain $S_{\rm ITL}^{\rm rec}(M)>S_{\rm ITL}^{\rm pred}(M)$, but we note that all numerical values ($\sim 10^{-3}$) are an order of magnitude smaller than for INs.
3.3 The effects of a mass dependence of matrix $\mathbf{C}$
We explored the effect on our method of a possible dependence of the confusion matrix on the mass of galaxies. For each mass bin considered in Fig. 2, we computed a confusion matrix, $\mathbf{C}(M)$, using only those galaxies in the test sample that have stellar mass in the bin. The matrices so obtained (not shown here) do have a mild dependence on mass. However, this dependence do not change the results discussed in the previous subsection in a significant way. In Fig. 3, we present the resulting values of $S_{i}^{\rm rec}(M)$ as orange lines, where it is evident that the differences compared to the case of a unique $\mathbf{C}$ are really small. We can safely affirm there is no need to use a mass-dependent (or absolute-magnitude-dependent) confusion matrix. Furthermore, a mass-dependent confusion matrix implies a greater dependence on the specifics of how the galaxy evolution model works. It is desirable to minimise reliance on the model as much as possible.
@@ 4 Application to an observed sample of X-ray clusters of galaxies
As an example of the potential of the proposed statistical technique to correct the effects of misclassification, we applied our method to the sample of X-ray clusters and Sloan Digital Sky Survey (SDSS) galaxies used in Martínez et al. (2023). The sample comprises 104 X-ray clusters drawn from the works by Coenda & Muriel (2009) and Muriel & Coenda (2014), where the authors compute a number of properties of clusters originally catalogued by Popesso et al. (2004) and Bohringer et al. (2000), respectively. Martínez et al. (2023) searched for all spectroscopic SDSS-DR7 (Abazajian et al. 2009) galaxies within cylindrical volumes in redshift space centred in the clusters, and with radius $3\times R_{200}$ in projection and a longitude of ±$\pm 3\times\sigma$ alongside the line-of-sight of the cluster. Galaxies were classified in the five classes under study using their positions in the PPSD, the roger-computed probabilities, and the thresholds of Coenda et al. (2022).
Using the confusion matrix computed in Sect. 3.2, we estimated the predicted and recovered distributions of the (${}^{0.1}(u-r)$ colour of the five classes in four bins in stellar mass. The resulting distributions are shown in Fig. 4. The outcome of applying our method to this sample of observed galaxies and clusters qualitatively agrees with our findings in Sect. 3.2, which were derived from applying the method to a simulated galaxy catalogue. We can observe basically the same trends of Fig. 3. For CL, the distributions recovered by our method eliminate most of the contribution from blue galaxies, this is more notorious in the lower-mass bins. According to the insights of Sect. 3.2, these blue galaxies are mostly actual RINs. In the case of BSs, the main feature we obtain is the removal of the contamination from blue galaxies, particularly at the lowest-mass bin. According to the confusion matrix we used (see Sect. 3.2), some $70\%$ of the contamination should be due to galaxies erroneously classified as INs. Regarding RINs, the result of our method is most notably seen in the removal of a population of red, almost exclusively CLs, most clearly in the lower-mass bins. As in Sect. 3.2, for INs, the method gives results very similar to the raw predictions. Finally, for ITLs, once again, there are virtually no differences between the predicted and recovered colour distributions.
@@ 5 Conclusions
We propose a method for recovering the distributions of observed properties of galaxies classified according to their position in the PPSD. The method requires knowledge of the confusion matrix, or at least an estimation of it, which is inverted and applied to observed distributions. We tested this method on simulated clusters and galaxies and applied it to a sample of galaxies and X-ray clusters from the SDSS. For both simulated and observed galaxies, we classified galaxies in the PPSD using the code roger from de los Rios et al. (2021) plus the criteria from Coenda et al. (2022). In particular, we utilised colour as the parameter for analysing the distributions. These two choices, namely (i) the classification using roger plus thresholds and (ii) colour, are not essential to our objectives. Any classification scheme over the PPSD, such as those by Rhee et al. (2017) and Pasquali et al. (2019), and any galaxy property would be suitable provided one has an estimation of the confusion matrix.
Using the inversion of the confusion matrix along with our roger-based classification provides a better estimation of the distribution of colours of galaxies in and around clusters. The method is successful at decontaminating different classes of galaxies from the contamination of other classes, and it is particularly useful for regions of the PPSD where the overlap between classes complicates the classification most. The proposed method enables further analyses of galaxies when the classification is an issue. It is straightforward to reconstruct some statistics of galaxy populations as a function of stellar mass (or absolute magnitude). For instance, from Fig. 4, we can compute the median of the colour distribution as a function of mass for the blue and red galaxy populations of each galaxy class.
We can draw some conclusions from the results we obtain for the five classes. The bulk of galaxies in clusters that are blue and less massive than $M\simeq 10^{10.5}h^{-1}M_{\odot}$ are RINs. In terms of colour, BSs are on average in between CLs and RINs, while being distinctly redder than INs. Thus, a single excursion through the inner regions of a cluster produces significant changes in galaxy colour, due to a quench in star formation (e.g. Hough et al. 2023; Ruiz et al. 2023).
@@ Appendix A Additional figure
Figure 4: Application of our method to the sample of galaxies in and around X-ray clusters used by Martínez et al. (2023). As in Fig. 2, rows show different classes, and columns different stellar mass ranges. Green shaded histograms are predicted colour distributions and violet shaded histograms recovered colour distributions.